OpenStudy (anonymous):

Tell me if this answer is correct The length of a rectangle is twice the width. The area is 162^2yds. Find the length and width Width: 9 Length: 18

OpenStudy (anonymous):

hey sai!

OpenStudy (anonymous):

Someone gave me this answer earlier, but I wanted a second opinion

OpenStudy (anonymous):

I double, and sometimes triple check my work

OpenStudy (saifoo.khan):

Hi Cali.

OpenStudy (saifoo.khan):

Let me seee,,,,

OpenStudy (saifoo.khan):

x(2x) = 162 2x^2 = 162 x^2 = 81 x^2 - 81 = 0 (x-9)(x+9) = 0 So, x = 9

OpenStudy (saifoo.khan):

Width = x = 9 Length = 2x = 18

OpenStudy (anonymous):

Thanks Sai!

OpenStudy (saifoo.khan):

No problem. It's Saif btw. LOL.

OpenStudy (anonymous):

sorry bout that Saif

OpenStudy (saifoo.khan):

No problem.